OpenStudy (anonymous):

Which system has (-8, 0) as the solution?

OpenStudy (nubeer):

is there any options?

OpenStudy (anonymous):

no sir

OpenStudy (nubeer):

ok.. i dnt know how to explain how i got the equation but it can be y =x +8

OpenStudy (anonymous):

x - 4y = -8 2x - 3y = -16